OpenStudy (anonymous):

Lori bought a 3-liter bottle of soda. Yesterday she drank 48 centiliters of the soda. How much of the soda was left?

OpenStudy (anonymous):

1 liter = 100 centiliter. So, 48 centiliter = 48/100 liter = 0.48 liter Amount of soda that remained = 3 - 0.48 liter = 2.52 liter